# Degree of Kripke-incompleteness of Tense Logics

### v1

@@ Abstract
The degree of Kripke-incompleteness of a logic $L$ in some lattice $\mathcal{L}$ of logics is the cardinality of logics in $\mathcal{L}$ which share the same class of Kripke-frames with $L$. A celebrated result on Kripke-incompleteness is Blok’s dichotomy theorem for the degree of Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K})$: every modal logic $L\in\mathsf{NExt}(\mathsf{K})$ is of the degree of Kripke-incompleteness $1$ or $2^{\aleph_{0}}$. In this work, we show that the dichotomy theorem for $\mathsf{NExt}(\mathsf{K})$ can be generalized to the lattices $\mathsf{NExt}(\mathsf{K}_{t})$, $\mathsf{NExt}(\mathsf{K4}_{t})$ and $\mathsf{NExt}({\mathsf{S4}_{t}})$ of tense logics. We also prove that in $\mathsf{NExt}(\mathsf{K}_{t})$, $\mathsf{NExt}(\mathsf{K4}_{t})$ and $\mathsf{NExt}({\mathsf{S4}_{t}})$, iterated splittings are exactly the strictly Kripke-complete logics.
@@ 0.1 Introduction
In general, studying the degree of Kripke-incompleteness in $\mathcal{L}$ amounts to analyzing the equivalence relation $\equiv_{\mathsf{Fr}}$ on $\mathcal{L}$, where $L_{1}\equiv_{\mathsf{Fr}}L_{2}$ iff $L_{1}$ shares the same class of frames as $L_{2}$, i.e., $\mathsf{Fr}(L_{1})=\mathsf{Fr}(L_{2})$. The degree of Kripke-incompleteness of $L$ is the cardinality of the equivalence class $[L]_{\equiv_{\mathsf{Fr}}}$ in $\mathcal{L}$. A logic $L$ is strictly Kripke-complete in $\mathcal{L}$ if $\mathsf{deg}_{\mathcal{L}}(L)=1$.
@@
A celebrated result in this field is Blok’s dichotomy theorem for the degree of Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K})$: every modal logic $L\in\mathsf{NExt}(\mathsf{K})$ is of the degree of Kripke-incompleteness $1$ or $2^{\aleph_{0}}$. This theorem was first proved in [3] algebraically by showing that union-splittings in $\mathsf{NExt}(\mathsf{K})$ are exactly the consistent strictly Kripke-complete logics and all other consistent logics have the degree $2^{\aleph_{0}}$. Blok’s characterization shows the connection between strictly Kripke-complete normal modal logics and splittings of lattices of logics in $\mathsf{NExt}(\mathsf{K})$. For more research on splittings of lattices of modal, tense and subframe logics, we refer the readers to [21, 14, 12, 18]. A proof based on relational semantics was given later in [4, Section 10.5]. The characterization of the degree of Kripke-incompleteness given by Blok indicates locations of Kripke-complete logics in the lattice $\mathsf{NExt}(\mathsf{K})$. Since Blok’s proof relies heavily on non-transitive frames, it is natural to ask whether the dichotomy theorem holds for sublattices of $\mathsf{NExt}(\mathsf{K})$, especially for the lattices of transitive modal logics such as $\mathsf{K4}$ and $\mathsf{S4}$. These problems remain open, see [4, Problem 10.5].
@@
Generally, one can always replace the class $\mathsf{Fr}$ of all Kripke frames with some proper class $\mathcal{C}$ of mathematical structures, for example, the class $\mathsf{MA}$ of all modal algebras, the class $\mathsf{NF}$ of all neighborhood frames and the class $\mathsf{Fin}$ of all finite frames. Let $\mathcal{L}=\mathsf{NExt}(\mathsf{K})$. Since every normal modal logic is complete with respect to modal algebras [4, Theorem 7.73] we obtain that $\equiv_{\mathsf{MA}}$ is the identity relation on $\mathcal{L}$. Note that logics that enjoy the finite model property (FMP) are Kripke complete and Kripke complete logics are neighborhood complete, we have $\equiv_{\mathsf{NF}}\subseteq\equiv_{\mathsf{Fr}}\subseteq\equiv_{\mathsf{Fin}}$. The degree of modal incompleteness with respect to neighborhood semantics was also well-investigated, e.g., by Chagrova [5], Dziobiak [10] and Litak [15]. Dziobiak [10] proved the dichotomy theorem for the degree of incompleteness in the lattice $\mathsf{Ext}(\mathsf{D}\oplus(\Box^{n}p\to\Box^{n+1}p))$ w.r.t neighborhood semantics for all $n\in\omega$. Litak [15] studied modal incompleteness w.r.t Boolean algebras with operators (BAOs) and showed the existence of a continuum of neighborhood-incomplete modal logics extending $\mathsf{Grz}$. For more on modal incompleteness from an algebraic view, we refer the readers to [16]. Bezhanishvili et al. [1] introduced the notion of the degree of FMP of $L$ in $\mathcal{L}$, which is in fact the cardinality of the equivalence class $[L]_{\equiv_{\mathsf{Fin}}}$. The anti-dichotomy theorem for the degree of FMP for extensions of the intuitionistic propositional logic $\mathsf{IPC}$ was proved in [1]: for each cardinal $\kappa$ with $0<\kappa\leq\aleph_{0}$ or $\kappa=2^{\aleph_{0}}$, there exists $L\in\mathsf{Ext}(\mathsf{IPC})$ such that the degree of FMP of $L$ in $\mathsf{Ext}(\mathsf{IPC})$ is $\kappa$. It was also shown in [1] that the anti-dichotomy theorem of the degree of FMP holds for $\mathsf{NExt}(\mathsf{K4})$ and $\mathsf{NExt}(\mathsf{S4})$. Degrees of FMP in lattices of bi-intuitionistic logics were studied in [9]. Given close connections between modal logics, bi-intuitionistic logics and tense logics, it is natural to study the degree of Kripke-incompleteness in lattices of tense logics.
@@
Tense logics are bi-modal logics that include a future-looking necessity modality $\Box$ and a past-looking possibility modality $\blacklozenge$, of which the lattices are substantially different from those of modal logics (see [13, 19, 17]). As far as we are aware, the degree of Kripke-incompleteness in lattices of tense logics has not been investigated systematically. In this work, we study Kripke-incompleteness in lattices of tense logics. We start with the lattice $\mathsf{NExt}(\mathsf{K}_{t})$ of all tense logics. Inspired by the proof for Blok’s dichotomy theorem in [4], we prove the dichotomy theorem for tense logics, that is, every tense logic $L\in\mathsf{NExt}(\mathsf{K}_{t})$ is of degree of Kripke-incompleteness $1$ or $2^{\aleph_{0}}$. This is proved by showing that union-splittings in $\mathsf{NExt}(\mathsf{K}_{t})$ are exactly the strictly Kripke-complete logics and all other logics have the degree $2^{\aleph_{0}}$. By a similar argument, we prove the dichotomy theorem of the degree of Kripke-incompleteness for $\mathsf{NExt}(\mathsf{K4}_{t})$. Finally, we turn to the lattice $\mathsf{NExt}({\mathsf{S4}_{t}})$. We provide the following characterization of the degree of Kripke-incompleteness in $\mathsf{NExt}({\mathsf{S4}_{t}})$: iterated splittings are strictly Kripke-complete and all other logics are of degree $2^{\aleph_{0}}$, where iterated splittings are intuitively splittings in the lattice of splitting logics. The dichotomy theorem of the degree of Kripke-incompleteness for $\mathsf{NExt}({\mathsf{S4}_{t}})$ follows from the characterization immediately. It also follows that in the lattice $\mathsf{NExt}({\mathsf{S4}_{t}})$, strictly Kripke-complete logics are no longer the union-splittings.
@@ 0.2 Preliminaries
The connectives ,¬$\top$,$\neg$,$\wedge$ , ¬ , ∧ and $\vee$ are defined as usual. Let $\Diamond\varphi:=\neg\Box\neg\varphi$ and $\blacksquare\varphi:=\neg\blacklozenge\neg\varphi$. The modal degree $md(\varphi)$ of a formula $\varphi$ is defined inductively as follows:
@@
A substitution is a homomorphism (.$(.)^{s}:\mathscr{L}_{t}\twoheadrightarrow\mathscr{L}_{t}$ on the formula algebra $\mathscr{L}_{t}$.
@@
A frame is a pair $\mathfrak{F}=(X,R)$ where $X$ is a nonempty set and $R\subseteq X\times X$. We write $Rxy$ when ${\langle x,y\rangle}\in R$. The inverse of $R$ is defined as $\breve{R}=\{{\langle y,x\rangle}:Rxy\}$. For every $x\in X$, let $R[x]=\{y\in X:Rxy\}$ and $\breve{R}[x]=\{u\in X:Ryx\}$. For every $U\subseteq X$, we define $R[U]=\bigcup_{x\in U}R[x]$ and $\breve{R}[U]=\bigcup_{x\in U}\breve{R}[x]$. Let $\mathsf{Fr}$ and $\mathsf{Fin}$ denote the class of all frames and finite frames, respectively.
A general frame is a triple $\mathbb{F}=(X,R,A)$ where $(X,R)$ is a frame and $A\subseteq\mathcal{P}(X)$ is a set such that $\varnothing\in A$ and $A$ is closed under the operators $\cap$, $(\cdot)^{c}$, $R[\cdot]$ and $\breve{R}[\cdot]$. We call $A$ the set of internal sets in $\mathbb{F}$. We write $\kappa\mathbb{F}$ for the underlying frame $(X,R)$ of $\mathbb{F}$. Let $\mathsf{GF}$ denote the class of all general frames.
@@
$V(\bot)$ ==$\varnothing$,= ∅ , $V(\varphi\to\psi)$ $=(V(\varphi))^{c}\cup V(\psi),$
@@
For each $n\in\mathbb{Z}^{+}$, we define the formulas $(\mathsf{alt}^{+}_{n})$, $(\mathsf{alt}^{-}_{n})$, $(\mathsf{bz}_{n})$, $(\mathsf{bw}^{+}_{n})$ and $(\mathsf{bw}^{-}_{n})$ as follows:
@@
Splittings of lattices of logics play a core role in Blok’s characterization of degree of Kripke-incompleteness. Consider the lattice $\mathsf{NExt}(\mathsf{K})$ of normal modal logics. It is known that a normal modal logic $L$ splits $\mathsf{NExt}(\mathsf{K})$ if and only if $L=\mathsf{Log}(\mathfrak{F})$ for some finite rooted cycle free frame $\mathfrak{F}$ (see [4, Theorems 10.49 and 10.53]). Blok [3] proved that a normal modal logic $L$ has the degree of incompleteness $1$ in $\mathsf{NExt}(\mathsf{K})$ if and only if $L$ is a union splitting of $\mathsf{NExt}(\mathsf{K})$. It was also showed in [3] that in the lattice $\mathsf{NExt}(\mathsf{K})$, consistent iterated splittings are exactly union-splittings.
@@
[…]
@@ 0.3 Reflective Unfolding of Kripke frames
($({\{{\langle w,0\rangle}\}}\times{\{{\langle x,1\rangle}:x\in S[u]\}})\cup({\{{\langle x,1\rangle}:x\in\breve{S}[u]\}}\times{\{{\langle w,0\rangle}\}})$
@@
  • for all $x,y\in X$, $T{\langle x,0\rangle}{\langle y,0\rangle}$ if and only if $Rxy$;
  • for all $x,y\in Y\setminus{\{u\}}$, $T{\langle x,1\rangle}{\langle y,1\rangle}$ if and only if $Sxy$;
  • for all $y\in Y\setminus{\{u\}}$, $T{\langle w,0\rangle}{\langle y,1\rangle}$ if and only if $Suy$;
  • for all $y\in Y\setminus{\{u\}}$, $T{\langle y,1\rangle}{\langle w,0\rangle}$ if and only if $Syu$.
@@ 0.4 Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K}_{t})$ and $\mathsf{NExt}(\mathsf{K4}_{t})$
In this section, we generalize Blok’s dichotomy theorem for $\mathsf{NExt}(\mathsf{K})$ to the lattices $\mathsf{NExt}(\mathsf{K}_{t})$ and $\mathsf{NExt}(\mathsf{K4}_{t})$ of tense logics. For $\mathsf{NExt}(\mathsf{K}_{t})$ and $\mathsf{NExt}(\mathsf{K4}_{t})$, we prove the dichotomy theorems by showing that the union-splittings of $\mathsf{NExt}(\mathsf{K}_{t})$ and $\mathsf{NExt}(\mathsf{K4}_{t})$ are exactly those having the degree of Kripke-incompleteness $1$, and all other tense logics have the degree of Kripke-incompleteness $2^{\aleph_{0}}$, respectively.

### v2

@@ Abstract
The degree of Kripke-incompleteness of a logic $L$ in some lattice $\mathcal{L}$ of logics is the cardinality of logics in $\mathcal{L}$ which share the same class of Kripke-frames with $L$. A celebrated result on Kripke-incompleteness is Blok’s dichotomy theorem for the degree of Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K})$: every modal logic $L\in\mathsf{NExt}(\mathsf{K})$ is of the degree of Kripke-incompleteness 11 or $2^{\aleph_{0}}$. In this work, we show that the dichotomy theorem for $\mathsf{NExt}(\mathsf{K})$ can be generalized to the lattices $\mathsf{NExt}(\mathsf{K}_{t})$, $\mathsf{NExt}(\mathsf{K4}_{t})$ and $\mathsf{NExt}({\mathsf{S4}_{t}})$ of tense logics. We also prove that in $\mathsf{NExt}(\mathsf{K}_{t})$, $\mathsf{NExt}(\mathsf{K4}_{t})$ and $\mathsf{NExt}({\mathsf{S4}_{t}})$, iterated splittings are exactly the strictly Kripke-complete logics.
@@ 0.1 Introduction
In general, studying the degree of Kripke-incompleteness in $\mathcal{L}$ amounts to analyzing the equivalence relation $\equiv_{\mathsf{Fr}}$ on $\mathcal{L}$, where $L_{1}\equiv_{\mathsf{Fr}}L_{2}$ iff $L_{1}$ shares the same class of frames as $L_{2}$, i.e., $\mathsf{Fr}(L_{1})=\mathsf{Fr}(L_{2})$. The degree of Kripke-incompleteness of $L$ is the cardinality of the equivalence class $L]_{\equiv_{\mathsf{Fr}}}$ in $\mathcal{L}$. A logic $L$ is strictly Kripke-complete in $\mathcal{L}$ if $\mathsf{deg}_{\mathcal{L}}(L)=1$.
@@
A celebrated result in this field is Blok’s dichotomy theorem for the degree of Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K})$: every modal logic $L\in\mathsf{NExt}(\mathsf{K})$ is of the degree of Kripke-incompleteness 11 or $2^{\aleph_{0}}$. This theorem was first proved in [3] algebraically by showing that union-splittings in $\mathsf{NExt}(\mathsf{K})$ are exactly the consistent strictly Kripke-complete logics and all other consistent logics have the degree $2^{\aleph_{0}}$. Blok’s characterization shows the connection between strictly Kripke-complete normal modal logics and splittings of lattices of logics in $\mathsf{NExt}(\mathsf{K})$. For more research on splittings of lattices of modal, tense and subframe logics, we refer the readers to [21, 14, 12, 18]. A proof based on relational semantics was given later in [4, Section 10.5]. The characterization of the degree of Kripke-incompleteness given by Blok indicates locations of Kripke-complete logics in the lattice $\mathsf{NExt}(\mathsf{K})$. Since Blok’s proof relies heavily on non-transitive frames, it is natural to ask whether the dichotomy theorem holds for sublattices of $\mathsf{NExt}(\mathsf{K})$, especially for the lattices of transitive modal logics such as $\mathsf{K4}$ and $\mathsf{S4}$. These problems remain open, see [4, Problem 10.5].
@@
Generally, one can always replace the class $\mathsf{Fr}$ of all Kripke frames with some proper class $\mathcal{C}$ of mathematical structures, for example, the class $\mathsf{MA}$ of all modal algebras, the class $\mathsf{NF}$ of all neighborhood frames and the class $\mathsf{Fin}$ of all finite frames. Let $\mathcal{L}=\mathsf{NExt}(\mathsf{K})$. Since every normal modal logic is complete with respect to modal algebras [4, Theorem 7.73] we obtain that $\equiv_{\mathsf{MA}}$ is the identity relation on $\mathcal{L}$. Note that logics that enjoy the finite model property (FMP) are Kripke complete and Kripke complete logics are neighborhood complete, we have $\equiv_{\mathsf{NF}}\subseteq\equiv_{\mathsf{Fr}}\subseteq\equiv_{\mathsf{Fin}}$. The degree of modal incompleteness with respect to neighborhood semantics was also well-investigated, e.g., by Chagrova [5], Dziobiak [10] and Litak [15]. Dziobiak [10] proved the dichotomy theorem for the degree of incompleteness in the lattice $\mathsf{Ext}(\mathsf{D}\oplus(\Box^{n}p\to\Box^{n+1}p))$ w.r.t neighborhood semantics for all $n\in\omega$. Litak [15] studied modal incompleteness w.r.t Boolean algebras with operators (BAOs) and showed the existence of a continuum of neighborhood-incomplete modal logics extending $\mathsf{Grz}$. For more on modal incompleteness from an algebraic view, we refer the readers to [16]. Bezhanishvili et al. [1] introduced the notion of the degree of FMP of $L$ in $\mathcal{L}$, which is in fact the cardinality of the equivalence class $L]_{\equiv_{\mathsf{Fin}}}$. The anti-dichotomy theorem for the degree of FMP for extensions of the intuitionistic propositional logic $\mathsf{IPC}$ was proved in [1]: for each cardinal $\kappa$ with $0<\kappa\leq\aleph_{0}$ or $\kappa=2^{\aleph_{0}}$, there exists $L\in\mathsf{Ext}(\mathsf{IPC})$ such that the degree of FMP of $L$ in $\mathsf{Ext}(\mathsf{IPC})$ is $\kappa$. It was also shown in [1] that the anti-dichotomy theorem of the degree of FMP holds for $\mathsf{NExt}(\mathsf{K4})$ and $\mathsf{NExt}(\mathsf{S4})$. Degrees of FMP in lattices of bi-intuitionistic logics were studied in [9]. Given close connections between modal logics, bi-intuitionistic logics and tense logics, it is natural to study the degree of Kripke-incompleteness in lattices of tense logics.
@@
Tense logics are bi-modal logics that include a future-looking necessity modality $\Box$ and a past-looking possibility modality $\blacklozenge$, of which the lattices are substantially different from those of modal logics (see [13, 19, 17]). As far as we are aware, the degree of Kripke-incompleteness in lattices of tense logics has not been investigated systematically. In this work, we study Kripke-incompleteness in lattices of tense logics. We start with the lattice $\mathsf{NExt}(\mathsf{K}_{t})$ of all tense logics. Inspired by the proof for Blok’s dichotomy theorem in [4], we prove the dichotomy theorem for tense logics, that is, every tense logic $L\in\mathsf{NExt}(\mathsf{K}_{t})$ is of degree of Kripke-incompleteness 11 or $2^{\aleph_{0}}$. This is proved by showing that union-splittings in $\mathsf{NExt}(\mathsf{K}_{t})$ are exactly the strictly Kripke-complete logics and all other logics have the degree $2^{\aleph_{0}}$. By a similar argument, we prove the dichotomy theorem of the degree of Kripke-incompleteness for $\mathsf{NExt}(\mathsf{K4}_{t})$. Finally, we turn to the lattice $\mathsf{NExt}({\mathsf{S4}_{t}})$. We provide the following characterization of the degree of Kripke-incompleteness in $\mathsf{NExt}({\mathsf{S4}_{t}})$: iterated splittings are strictly Kripke-complete and all other logics are of degree $2^{\aleph_{0}}$, where iterated splittings are intuitively splittings in the lattice of splitting logics. The dichotomy theorem of the degree of Kripke-incompleteness for $\mathsf{NExt}({\mathsf{S4}_{t}})$ follows from the characterization immediately. It also follows that in the lattice $\mathsf{NExt}({\mathsf{S4}_{t}})$, strictly Kripke-complete logics are no longer the union-splittings.
@@ 0.2 Preliminaries
The connectives $\top,\neg,\wedge$ and $\vee$ are defined as usual. Let $\Diamond\varphi:=\neg\Box\neg\varphi$ and $\blacksquare\varphi:=\neg\blacklozenge\neg\varphi$. The modal degree $md(\varphi)$ of a formula $\varphi$ is defined inductively as follows:
@@
A substitution is a homomorphism $(.)^{s}:\mathscr{L}_{t}\twoheadrightarrow\mathscr{L}_{t}$ on the formula algebra $\mathscr{L}_{t}$.
@@
A frame is a pair $\mathfrak{F}=(X,R)$ where $X$ is a nonempty set and $R\subseteq X\times X$. We write RxyRxy when ${\langle x,y\rangle}\in R$. The inverse of $R$ is defined as $\breve{R}=\{{\langle y,x\rangle}:Rxy\}$. For every $x\in X$, let $R[x]=\{y\in X:Rxy\}$ and $\breve{R}[x]=\{u\in X:Ryx\}$. For every $U\subseteq X$, we define $R[U]=\bigcup_{x\in U}R[x]$ and $\breve{R}[U]=\bigcup_{x\in U}\breve{R}[x]$. Let $\mathsf{Fr}$ and $\mathsf{Fin}$ denote the class of all frames and finite frames, respectively.
A general frame is a triple $\mathbb{F}=(X,R,A)$ where $(X,R)$ is a frame and $A\subseteq\mathcal{P}(X)$ is a set such that $\varnothing\in A$ and AA is closed under the operators $\cap$, $(\cdot)^{c}$, $R[\cdot]$ and $\breve{R}[\cdot]$. We call AA the set of internal sets in $\mathbb{F}$. We write $\kappa\mathbb{F}$ for the underlying frame $(X,R)$ of $\mathbb{F}$. Let $\mathsf{GF}$ denote the class of all general frames.
@@
$V(\bot)$ $=\varnothing,$ $V(\varphi\to\psi)$ $=(V(\varphi))^{c}\cup V(\psi),$
@@
[…]
@@
Splittings of lattices of logics play a core role in Blok’s characterization of degree of Kripke-incompleteness. Consider the lattice $\mathsf{NExt}(\mathsf{K})$ of normal modal logics. It is known that a normal modal logic $L$ splits $\mathsf{NExt}(\mathsf{K})$ if and only if $L=\mathsf{Log}(\mathfrak{F})$ for some finite rooted cycle free frame $\mathfrak{F}$ (see [4, Theorems 10.49 and 10.53]). Blok [3] proved that a normal modal logic $L$ has the degree of incompleteness 11 in $\mathsf{NExt}(\mathsf{K})$ if and only if $L$ is a union splitting of $\mathsf{NExt}(\mathsf{K})$. It was also showed in [3] that in the lattice $\mathsf{NExt}(\mathsf{K})$, consistent iterated splittings are exactly union-splittings.
@@
Lattices of tense logics are substantially different from those of modal logics. Splittings of lattices of tense logics have been investigated in [13]. Let $\bullet$ denote the frame $({\{0\}},\varnothing)$. Then
@@ 0.3 Reflective Unfolding of Kripke frames
$({\{{\langle w,0\rangle}\}}\times{\{{\langle x,1\rangle}:x\in S[u]\}})\cup({\{{\langle x,1\rangle}:x\in\breve{S}[u]\}}\times{\{{\langle w,0\rangle}\}})$
@@
  • for all $x,y\in X$, $T{\langle x,0\rangle}{\langle y,0\rangle}$ if and only if RxyRxy;
  • for all $x,y\in Y\setminus{\{u\}}$, $T{\langle x,1\rangle}{\langle y,1\rangle}$ if and only if SxySxy;
  • for all $y\in Y\setminus{\{u\}}$, $T{\langle w,0\rangle}{\langle y,1\rangle}$ if and only if SuySuy;
  • for all $y\in Y\setminus{\{u\}}$, $T{\langle y,1\rangle}{\langle w,0\rangle}$ if and only if SyuSyu.
@@ 0.4 Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K}_{t})$ and $\mathsf{NExt}(\mathsf{K4}_{t})$
In this section, we generalize Blok’s dichotomy theorem for $\mathsf{NExt}(\mathsf{K})$ to the lattices $\mathsf{NExt}(\mathsf{K}_{t})$ and $\mathsf{NExt}(\mathsf{K4}_{t})$ of tense logics. For $\mathsf{NExt}(\mathsf{K}_{t})$ and $\mathsf{NExt}(\mathsf{K4}_{t})$, we prove the dichotomy theorems by showing that the union-splittings of $\mathsf{NExt}(\mathsf{K}_{t})$ and $\mathsf{NExt}(\mathsf{K4}_{t})$ are exactly those having the degree of Kripke-incompleteness 11, and all other tense logics have the degree of Kripke-incompleteness $2^{\aleph_{0}}$, respectively.